If Leah is 6 years older than Sue, and John is 5 years older than Leah, and the total of their ages is 41, then how old is Sue?
Answer by mananth(16946)

Let sue's age be x
Leah will be x+6
John will be x+6+5
...
Their sum = 41
x+x+6+x+6+5=41
3x+17=41
-17 -17
3x=41-17
3x=24
/3
3x/3 = 24/3
x=8 Sue 's age
Leah is x+6 = 8+6 = 14
John = x+6+5 = 8+6+5 = 19
...
CHECK
8+14+19=41
...
m.ananth@hotmail.ca